# From Estimands to Robust Inference of Treatment Effects in Platform Trials

### v1

@@ 3.1 Randomization procedures in platform trials
Consider a platform trial with $J$ treatments and $n$ patients. For each patient $i$, let $\bm{W}_{i}$ be a vector of all observed baseline variables (e.g., baseline covariates and enrollment window by calendar time) not affected by the treatment, $A_{i}$ be the treatment assignment indicator that equals $j$ if the patient is assigned to arm $j$, $Y_{i}^{(j)}$ be the potential outcome under treatment $j$, $j=1,...,J$ (Neyman,, 1923; Rubin,, 1974), and $Y_{i}=Y_{i}^{(A_{i})}$ be the observed outcome. Throughout, $(\bm{W}_{i},Y_{i}^{(1)},\dots,Y_{i}^{(J)},A_{i})$, $i=1,\dots,n$, are independent and identically distributed realizations of $(\bm{W},Y^{(1)},\dots,Y^{(J)},A)$ with finite second-order moments. The following assumption describes the randomization procedure in platform trials.
Assumption 1 (Randomization).
@@
There exists an observed baseline discrete variable ${Z}$ such that $A\perp(\bm{W},Y^{(1)},\dots,Y^{(J)})\mid{Z}$ and $P(A=j\mid{Z})=\pi_{j}({Z}),$ where $\perp$ stands for independence, $\mid$ stands for conditioning, $10\leq\pi_{j}({Z})<$ 1, $j=1,...,J$, are known treatment assignment probabilities once ${Z}$ is known, and $\sum_{j=1}^{J}\pi_{j}({Z})=1$.
@@
Assumption 1 highlights a fundamental difference between the randomization procedure used in platform trials versus traditional clinical trials. In traditional trials, the randomization probability is typically a fixed positive constant. However, in platform trials, this probability, $\pi_{j}(Z)$, varies depending on a baseline discrete variable $Z$ and may be zero for some treatments. Notably, $Z$ usually includes the enrollment window by calendar time and some key patient characteristics, as the enrollment window determines which treatments are available for enrollment, and patient characteristics determine the treatments for which a patient is eligible. Hence, $\pi_{j}(Z)$ may be zero if treatment $j$ has not entered the platform trial or has closed its enrollment, or if the patient does not meet the eligibility requirements for treatment $j$. Additionally, non-zero values of $\pi_{j}(Z)$ can differ over time and patient characteristics. Values of $\pi_{j}({Z})$, $j=1,...,J$, are known by design, but $\pi_{j}({Z})$ should be viewed as random due to the randomness of ${Z}$.
The example in Figure 1 illustrates these dynamics. Here, $Z=(Z_{\rm win},Z_{\rm sub},Z_{\rm B+})$ represent the enrollment window by calendar time, disease subtype (A or not A), and biomarker B (positive or negative), respectively, where $Z_{\rm B+}$ does not affect treatment assignment until enrollment window 5. For simplicity, consider randomization with equal probabilities. If patient $i$ with subtype A enters the trial during enrollment window 1 in which only sub-studies 1 and 2 are open, this patient is randomized to either sub-study with equal probability. If randomized to sub-study 1, the patient is further randomized to treatment 1 or 2 with equal probability. If randomized to sub-study 2, the patient is further randomized to treatment 1 or 3 with equal probability. Consequently, for this patient $i$, $\pi_{1}({Z}_{i})=0.5\times 0.5+0.5\times 0.5=0.5$, $\pi_{2}({Z}_{i})=0.5\times 0.5=0.25$, $\pi_{3}({Z}_{i})=0.5\times 0.5=0.25$, with all other $\pi_{j}({Z}_{i})$ values equal to 0. If another patient $i^{\prime}$ without subtype A enters the trial during the same enrollment window 1, then this patient is assigned to sub-study 1 with probability 1. Consequently, $\pi_{1}({Z}_{i^{\prime}})=\pi_{2}({Z}_{i^{\prime}})=0.5$ and all other $\pi_{j}({Z}_{i^{\prime}})$ values are 0. Similarly, if another patient $i{{}^{\prime\prime}}$ with subtype A enters the trial during enrollment window 2 in which sub-studies 1,2,3 are open, this patient has $\pi_{1}(Z_{i^{\prime\prime}})=0.5$, $\pi_{2}(Z_{i^{\prime\prime}})=\pi_{3}(Z_{i^{\prime\prime}})=\pi_{4}(Z_{i^{\prime\prime}})=1/6$, with all other $\pi_{j}({Z}_{i^{\prime\prime}})$ values equal to 0. These cases exemplify how the positivity and values of $\pi_{j}(Z)$ can vary with $Z$.
3.2 Treatment effects for concurrently eligible patients
@@
We first define a clinically meaningful estimand (ICH E9 , 2019, R1) for comparing a given pair of treatments, $j$ and $k$, one of which is typically a control. As discussed in the introduction, many researchers and the FDA have recommended that the comparison be based on concurrently eligible patients, i.e., patients who meet the eligibility criteria for both treatments and are enrolled during a time period when both treatments are available, and therefore could have been randomized to either treatment $j$ or $k$. Formally, the concurrently eligible patients for treatments $j$ and $k$ are those with $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$.
We provide two examples to illustrate this concept. The first example is based on Figure 1. When an investigator compares treatments 1 and 3 at the end of sub-study 2, the concurrently eligible patients are all subtype A patients from enrollment windows 1 and 2, which includes all patients in sub-study 2, as well as patients with subtype A in sub-study 1 and all patients in sub-study 3 from the first two windows. Patients without subtype A or enrolled in later windows are not included, as they are either ineligible or nonconcurrent. The second example is based on the SIMPLIFY design in Figure 2. The original analysis by Mayer-Hamblett et al., (2023) calculated the effects of discontinuing vs. continuing HS or DA within each study separately. However, this approach did not utilize all concurrently eligible patients with additional control data that could have been used from participants assigned to continue therapy across either study. Thus, when comparing discontinuing HS versus continuing therapy, for instance, the concurrently eligible patients are all individuals on HS regardless of their DA status, which includes all participants in the HS study, as well as those on both HS and DA in the DA study.
@@
We propose the following estimand to define the treatment effect for two given treatments $j$ and $k$ among concurrently eligible patients:
@@
where both subscripts $j$ and $k$ are used in $\bm{\vartheta}_{jk}$, $\theta_{jk}$, and $\theta_{kj}$ because both $\pi_{j}({Z})$ and $\pi_{k}({Z})$ are involved. In essence, a contrast of the estimand in (1) is an unconditional treatment effect for all concurrently eligible patients, averaging across all eligible patients and over time during which the enrollment window is open. Time can represent many background factors that may be changing although this is not explicitly recognized in estimand specification for traditional clinical trials.
Our defined population of concurrently eligible patients is the largest population that can preserve the integrity of randomization when inferring treatment effects. Individuals outside this group have zero probability of being assigned to treatments $j$ or $k$, or both, due to their ineligibility or the unavailability of treatments, thereby necessitating extrapolation. This is the reason why we do not consider the population means $E(Y^{(j)})$ and $E(Y^{(k)})$ that are defined for the entire platform trial population. Conversely, our estimand in (1) represents the treatment effect for a broader population compared to just those enrolled in a specific sub-study. Our estimand in (1) truly respects the eligibility criteria for the given pair of treatments for comparison, rather than on the subset assigned to a particular sub-study.
Throughout the paper, we use the notation convention that $a_{kj}$ is defined by swapping subscripts $j$ and $k$ in $a_{jk}$, e.g., $\theta_{jk}$ and $\theta_{kj}$ defined in (1). Based on the estimand ${\bm{\vartheta}}_{jk}$ in (1), the treatment effect of $j$ versus $k$ can be defined as the linear contrast $\theta_{jk}-\theta_{kj}$, the ratio $\theta_{jk}/\theta_{kj}$, or some other function of ${\bm{\vartheta}}_{jk}$. Our approach to defining treatment effects is adaptable to situations where population characteristics other than means are considered, or when the treatment effect for a targeted subgroup is of primary interest.
@@ 4.1 Naive method
After we define the estimand ${\bm{\vartheta}}_{jk}$ in (1), the next challenge is estimating each component of ${\bm{\vartheta}}_{jk}$ in (1), because even among the concurrently eligible patients defined by $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$, these non-zero probabilities $\pi_{j}({Z}),\pi_{k}({Z})$ may still vary over time and patient characteristics, leading to confounding.
A naive method is to estimate the component $\theta_{jk}$ by the sample mean $\bar{Y}_{jk}$ of $Y_{i}$’s from patients with $\pi_{j}({Z}_{i})>0$, $\pi_{k}({Z}_{i})>0$, and $A_{i}=j$. While $\bar{Y}_{jk}$ is valid in traditional trials, it is typically biased under platform trials as it overlooks confounding by ${Z}$, i.e., the fact that $Z$ affects both treatment and outcome. Specifically, as shown in the Supplementary Material, as $n\to\infty$,
@@
The limit of $\bar{Y}_{jk}$ in (2) is not equal to $\theta_{jk}$ in (1) unless either $\pi_{j}({Z})$ does not vary with ${Z}$ or $Y^{(j)}\perp{Z}$. Take the trial described in Section 2.1 and Figure 1 as an example, where $\pi_{j}({Z})$ varies with ${Z}$, the condition $Y^{(j)}\perp{Z}$ unrealistically rules out the influence of disease subtype on the outcome, as well as potential temporal effects, such as later recruited patients being healthier than those recruited earlier. The problem with the naive estimator $\bar{Y}_{jk}$ arises from confounding by $\pi_{j}({Z})$, not from conditioning on $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$. In traditional clinical trials, $\bar{Y}_{jk}$ is valid because $\pi_{j}({Z})$ does not vary with ${Z}$.
@@ 4.2 Inverse probability weighting
As noted in Section 4.1, the naive method is biased because it fails to account for confounding by $Z$. Under Assumption 1, since the true probability of receiving treatment $\pi_{j}(Z)$ is known, this confounding can be addressed using inverse probability weighting (IPW). The identification formula is (the proof is in the Supplementary Material):
$\theta_{jk}=E\bigg{\{}\frac{I(A=j)Y}{\pi_{j}({Z})}\ \Big{|}\ \pi_{j}({Z})>0,\pi_{k}({Z})>0\bigg{\}},\vspace{-3mm}$ (3)
@@
$\hat{\theta}_{jk}^{\rm\,ipw}=\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)Y_{i}}{\pi_{j}({Z}_{i})},\vspace{-3mm}$
@@
This IPW estimator may be improved by the following stabilized IPW (SIPW) (Hájek,, 1971; Robins et al.,, 2000) estimator, which normalizes the weights $I(A_{i}=j)/\pi_{j}({Z}_{i})$ for $i\in{{{\cal I}}}_{jk}$ to sum to 1:
@@
The IPW and SIPW methods utilize some covariate information through weighting with $\pi_{j}({Z})$. To improve efficiency, we can adjust for ${Z}$ and other baseline covariates using a model-assisted approach (Tsiatis et al.,, 2008; Ye et al.,, 2023; Bannick et al.,, 2023), which incorporates these covariates through a working model between the outcome and covariates and produces an asymptotically unbiased estimator even if the working model is incorrect. This robustness makes model-assisted approaches increasingly popular and recommended by the regulatory agencies (ICH E9,, 1998; EMA,, 2015; FDA,, 2021).
Specifically, let $\bm{X}\subset\bm{W}$ denote the observed vector of baseline covariates for adjustment, which may contain or overlap with ${Z}$. We estimate $E(Y^{(j)}\mid\bm{X},\pi_{j}({Z})>0,\pi_{k}({Z})>0,A=j)$ by $\hat{\mu}_{jk}(\bm{X})$, using a potentially misspecified working model. For continuous outcomes, $\hat{\mu}_{jk}(\bm{X})$ is usually from least squares regression of $Y_{i}$ on $\bm{X}$ using data where $\pi_{j}({Z})>0,\pi_{k}({Z})>0$ and $A=j$. For binary outcomes, $\hat{\mu}_{jk}(\bm{X})$ is often obtained from a logistic working model. After $\hat{\mu}_{jk}(\cdot)$ is obtained, we can adjust for covariate $\bm{X}$ using the augmented inverse probability weighting (AIPW) (Robins et al.,, 1994), resulting in the following AIPW estimator of $\theta_{jk}$,
@@
Interestingly, in its last term, the AIPW estimator leverages covariate information from all concurrently eligible patients in ${{{\cal I}}}_{jk}=\{i:\pi_{j}({Z}_{i})>0,\pi_{k}({Z}_{i})>0\}$, including those not receiving treatments $j$ or $k$. For example, in Figure 1, when estimating the effect of treatment 3 versus treatment 1 at the end of enrollment window 2 using AIPW, the estimator also incorporates covariate data from subtype A patients assigned to treatment 2 (in sub-study 1) and treatment 4 (in sub-study 3) during enrollment windows 1 and 2. These covariate data are not used in the IPW and SIPW estimators. Thus, AIPW offers a robust method to pool data from all concurrently eligible patients, thereby further extracting the potential of efficiency gain through platform trials.
@@
Both AIPW and SAIPW estimators are asymptotically normal (see Theorem 1 in Section 5). As discussed in Section 5, whether to stabilize the weights in the AIPW approach often results in asymptotically equivalent estimators.

### v4

@@ 3.1 Randomization procedure
Consider a platform trial with $J$ treatments. For each individual, let $\bm{W}$ be a vector of all observed baseline variables (e.g., baseline covariates and enrollment window by calendar time) not affected by the treatment, $A$ be the treatment assignment indicator that equals $j$ if the individual is assigned to arm $j$, $Y^{(j)}$ be the potential outcome under treatment $j$, $j=1,...,J$ (Neyman,, 1923; Rubin,, 1974), and $Y=Y^{(A)}$ be the observed outcome. We assume throughout that $(\bm{W}_{i},Y_{i}^{(1)},\dots,Y_{i}^{(J)},A_{i})$, $i=1,\dots,n$, is an independent and identically distributed sample of size $n$ from $(\bm{W},Y^{(1)},\dots,Y^{(J)},A)$ with finite second-order moments. The following is assumed for randomization in platform trials.
Assumption 1 (Constrained and non-uniform randomization).
@@
There exists an observed baseline variable ${Z}$ such that $A\perp(\bm{W},Y^{(1)},\dots,Y^{(J)})\mid{Z}$ and $P(A=j\mid{Z})=\pi_{j}({Z}),$ where $\perp$ stands for independence, $\mid$ stands for conditioning, $10\leq\pi_{j}({Z})<$ 1, $j=1,...,J$, are known treatment assignment probabilities once ${Z}$ is known, and $\sum_{j=1}^{J}\pi_{j}({Z})=1$.
@@
Assumption 1 highlights a fundamental difference between platform and traditional trials. Traditional trials typically use a fixed randomization probability, whereas in platform trials, the randomization probabilities, $\pi_{j}(Z)$, vary with a baseline variable $Z$, which usually is discrete and may include enrollment window (defined by calendar time), site, and individual characteristics. $\pi_{j}(Z)$ may even be zero for some treatments. For example, if treatment $j$ is not yet available, has closed enrollment, is not offered at a given site, or the individual is ineligible. Even when non-zero, $\pi_{j}(Z)$ can vary across individuals and over time. Although $\pi_{j}({Z})$ is known by design, it is random due to the randomness of ${Z}$.
[…]
3.2 Population
Building on the discussion in Section 1.2, we now formally define the population of interest for comparing two treatments $j$ and $k$ (one of which is typically a control). In this paper, population specifically refers to the population of individuals used to define the estimand.
Definition (ECE population).
@@ 3.2 Population
For the given treatments $j$ and $k$, the entire concurrently eligible (ECE) population is a population of all individuals who meet the eligibility criteria for both treatments and could potentially be enrolled during a time period when both treatments are available, and therefore could have been randomized to either treatment. Formally, the ECE population is a population of all individuals with $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$.
The ECE population is a population of all individuals who could potentially be assigned to treatments $j$ and $k$, rather than the population represented by those actually assigned to these treatments (i.e., the population represented by $A\in\{j,k\}$). In traditional trials, this distinction is often overlooked because the potential and actual populations coincide. However, in platform trials, this issue has been brought to the forefront.
We present two basic cases to illustrate why using the population represented by those actually assigned to the two treatments or sub-studies may lead to unexpected problems. In Case I, all participants are eligible for all treatments, and the focus is on how changing treatment allocation probabilities over time (non-uniform treatment assignment) can cause issues in defining the population. In Case II, all treatments enroll concurrently, and the focus is on how different eligibility criteria for different treatments (constrained treatment assignment) complicates population definition. Finally, we revisit the SIMPLIFY trial, which features both constrained and non-uniform treatment assignment, reflected in time-varying treatment allocation probabilities and varying eligibility criteria across treatments.
Figure 3: Comparison of traditional parallel-group and platform trial designs in a hypothetical mild to moderate disease population (Case I). In this case, the ECE population for comparing treatments 2 and 1 coincides with the entire trial population across all designs. In contrast, in platform trials, the population represented by those assigned to specific treatments or sub-studies may vary arbitrarily depending on the trial design or randomization probabilities.
Case I in Figure 3 shows various trial designs evaluating treatments 1 and 2, with all participants eligible for all treatments. The entire trial population is 50% mild and 50% moderate, but patient characteristics shift over time: enrollment window 1 representing one third of the population consists of 30% mild and 70% moderate, while enrollment window 2 comprising the remaining two third includes 60% mild and 40% moderate. We focus on the population for comparing treatments 2 and 1, as these two treatments are open for enrollment during the same period across all designs.
If a traditional parallel-group trial design (Design 1) is used, all participants have a positive probability of receiving either treatment, making the ECE population identical to the entire trial population (50% mild and 50% moderate). This is typically the default population for comparing treatments 1 and 2. The same applies to all platform trial designs in Figure 3, where the ECE population remains identical to the entire trial population regardless of the specific platform trial type (Design 2 vs. 3) or randomization probabilities (Design 2 vs. 2’). This consistency provides a principled justification for using the ECE population, as the target population for treatment comparisons should not depend on the specific trial operation formate or randomization probabilities, in the same way that choices of randomization ratio do not affect the study population in traditional trials.
However, in platform trials, it is common to conduct analyses that naively subset participants assigned to treatment 1 or 2 (Design 2 and 2’), or to sub-study 1 (Design 3). Even without explicit subsetting, methods such as the t-test are often applied in Designs 2 and 2’, where the randomization ratio between treatments 1 and 2 is constant. These analyses target the population of individuals assigned to specific treatments or sub-studies. Although such analysis populations still consist of concurrently eligible individuals, as shown in Figure 3, they differ from the ECE population and can be unexpectedly influenced by design formats and randomization ratios. See Section S1 of the Supplement for more discussions.
Lastly, we remark that the practice of focusing on the analysis populations assigned to specific treatments or sub-studies likely originates from traditional parallel-group trials. In traditional trials, these issues do not arise because the constant randomization ratio ensures that the analysis population in any arm coincides with the ECE population.
Figure 4: A hypothetical umbrella trial (Case II) with a population with different disease severity (mild, moderate, and severe). In this case, the ECE population for comparing treatments 2 and 1 is the mild-to-moderate disease population (with 50% mild and 50% moderate). In contrast, the analysis population defined by participants assigned to sub-study 1 may vary arbitrarily depending on how individuals with moderate disease are randomized between the two sub-studies.
Case II considers an umbrella trial where all treatments enroll concurrently. There are two sub-studies: one evaluates treatments 1 and 2 in a population with mild to moderate disease, and the other evaluates treatments 3 and 4 in a population with moderate to severe disease. The moderate disease population is randomized between the two sub-studies with a 1:1 ratio. These two populations are commonly of interest in the context of ulcerative colitis (FDA,, 2022). Suppose the overall disease population consists of 40% mild, 40% moderate, and 20% severe disease.
When studying treatments for mild-to-moderate disease, the target population is 50% mild and 50% moderate. As shown in Figure 4, the ECE population for comparing treatments 1 and 2 includes all individuals with mild or moderate disease, since all can potentially be assigned to these two treatments, which exactly represents the population of interest – 50% mild and 50% moderate. In contrast, if only considering those participants assigned to treatments 1 or 2, or those within sub-study 1, the analysis population may arbitrarily depend on the randomization ratio between sub-studies. For instance, if the allocation ratio to sub-study 1 for the moderate disease population decreases from 1/2 to 1/4, the resulting population restricted to sub-study 1 shifts to 80% mild and 20% moderate.
The SIMPLIFY trial has two sub-studies: one compares treatments 1 and 2 in a population on HS only or both HS and DA; the other compares treatments 1 and 3 in a population on DA only or both HS and DA. Participants on both HS and DA are first randomized between the two sub-studies using a 1:1 ratio, which was later adjusted to 3:1. As shown in Figure 1, the ECE population for comparing treatments 1 and 2 (i.e., assessing the effect of discontinuing HS versus continuing) includes all individuals on HS only or on both HS and DA, irrespective of the sub-study randomization ratio. This appropriately reflects the eligibility criteria for the treatments being compared. In contrast, a naive analysis restricted to participants assigned to the HS study targets an analysis population that differentially weights the HS-only and HS+DA groups, where the weights are determined by the randomization ratio among individuals on both HS and DA, as well as the timing of any changes to that ratio; further discussions are in section 6. Alternatively, an analysis that includes all participants assigned to treatments 1 or 2 may include DA-only individuals, making the resulting analysis population less relevant for evaluating the effect of discontinuing HS.
In summary, the ECE population preserves the integrity of randomized comparison, is invariant to the randomization ratio or trial type, and aligns with the intention-to-treat principle (ICH E9,, 1998). It is an appropriate population for both traditional and platform trials, and is especially useful for providing a clear and logical framework for defining populations when evaluating treatment effects in platform trials. In contrast, any population defined by actual treatment assignments or sub-study allocations may not be reasonable, as they can be quite arbitrary and influenced by trial design or specific randomization probabilities. Additionally, individuals outside the ECE population have zero probability of being assigned to either treatment $j$ or $k$, due to ineligibility or unavailability of treatments, thereby necessitating extrapolation.
In some situations, other populations, such as subgroups defined by specific characteristics, enrollment windows, or re-weighted populations from different enrollment windows, may be of interest. For example, in COVID-19 trials with rapidly evolving variants, the population enrolled during later windows, when newer variants are prevalent, may hold the greatest clinical relevance. The ECE population weights each enrollment window according to its natural size, but alternative weights that reflect clinical priorities can also be considered. The concept of ECE population can also be extended to comparisons involving more than two arms. For instance, the ECE population for comparing three treatments $j$, $k$, and $l$ is a population of all individuals for whom $\pi_{j}(Z)>0$, $\pi_{k}(Z)>0$, and $\pi_{l}(Z)>0$. In any case, we recommend starting with the ECE population as the baseline, then considering its subgroups or applying weighting to reflect the target population of interest. This approach offers clarity and specificity in defining the population of interest, ensuring that the population reflects the clinical question rather than having arbitrary dependence on design choices like randomization ratio or trial types. Moreover, the ECE population is crucial in other areas of statistical research, such as when using nonconcurrent controls to improve trial efficiency, as it provides a clear reference point for evaluating both efficiency gains and potential bias. These topics will be explored in future research.
3.3 Treatment effect for the ECE population
@@
With the population now clearly defined, we propose the following estimand to define the treatment effect for two given treatments $j$ and $k$ within the ECE population:
@@
where $\theta_{jk}$ and $\theta_{kj}$ depend on $j$ and $k$ because the ECE population depends on them. A contrast of the estimand in (1) is an unconditional treatment effect for the ECE population, averaging across all eligible participants and over time during which the enrollment window is open. For example, the treatment effect of $j$ versus $k$ can be defined as the linear contrast $\theta_{jk}-\theta_{kj}$, the ratio $\theta_{jk}/\theta_{kj}$, or some other function of ${\bm{\vartheta}}_{jk}$. In the context of the SIMPLIFY trial, $\theta_{21}-\theta_{12}$ is the effect of discontinuing HS versus continuing in a population of all individuals taking only HS or both HS and DA. A similar interpretation applies to $\theta_{31}-\theta_{13}$. This estimand framework is readily extendable to settings with multiple treatment arms.
Due to differences in the populations, $\theta_{jk}$ and $\theta_{kj}$ in (1) differ from $E(Y^{(j)})$ and $E(Y^{(k)})$, respectively, which are the means for the entire master protocol trial population and may include individuals who are non-concurrent or ineligible for treatments $j$ and $k$. Inference on $E(Y^{(j)})$ and $E(Y^{(k)})$ requires extrapolation. When these quantities are clinically of interest, one can apply methods to generalize the estimand in (1) to the entire trial population under additional assumptions (Degtiar and Rose,, 2023). Additionally, the estimand in (1) differs from the treatment effect for those assigned to specific arms (e.g., $E\{Y^{(j)}-Y^{(k)}\mid A\in(j,k)\}$), which, as shown by the formula, can depend on the distribution of $A$ and have problems such as arbitrary dependence on the randomization ratio. Similarly, the estimand in (1) also differs from the treatment effect for those assigned to specific sub-studies.
@@ 4 Estimation Methods
4.1 Analysis set
The analysis set, distinct from the population used to define the estimand, refers to the sample used in statistical analysis. While the ECE population is uniquely specified, the analysis set is not unique and can include multiple options. To maximize efficiency, the analysis set can include all concurrently eligible individuals in the index set ${{{\cal I}}}_{jk}=\{i:\pi_{j}({Z}_{i})>0,\pi_{k}({Z}_{i})>0\}$. However, practical considerations may lead to using a subset of ${{{\cal I}}}_{jk}$. For example, when applying a partial blinding strategy (FDA, 2023b, ) or if other sub-studies remain blinded at the time of analysis, it may be preferable to restrict the analysis set to individuals within a specific sub-study. In these cases, the analysis set is a biased sample of the ECE population and requires appropriate corrections. A general discussion of analysis sets and the conditions they must satisfy is provided in Section S1 of the Supplement. To focus our discussions, in the following, we consider either the analysis set defined by ${{{\cal I}}}_{jk}$ (with covariate adjustment) or the subset ${{{\cal I}}}_{jk}\cap\{i:A_{i}=j\text{ or }A_{i}=k\}$ (without covariate adjustment), where $\cap$ denotes the intersection of two sets.
@@ 4.2 Naive method
After we define the estimand, the next challenge is estimating each component of ${\bm{\vartheta}}_{jk}$ in (1), because even among the concurrently eligible individuals defined by $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$, these non-zero probabilities $\pi_{j}({Z})$ and $\pi_{k}({Z})$ may still vary over time and individual characteristics, leading to confounding.
A naive method is to estimate the component $\theta_{jk}$ by the sample mean $\bar{Y}_{jk}$ of $Y_{i}$’s from individuals with $\pi_{j}({Z}_{i})>0$, $\pi_{k}({Z}_{i})>0$, and $A_{i}=j$. However, $\bar{Y}_{jk}$ is typically biased for $\theta_{jk}$ under platform trials as it overlooks confounding by ${Z}$, i.e., the fact that $Z$ affects both treatment and outcome. Specifically, as shown in the Supplement, as $n\to\infty$,
@@
The limit of $\bar{Y}_{jk}$ in (2) is not equal to $\theta_{jk}$ in (1) unless either $\pi_{j}({Z})$ does not vary with ${Z}$ or $Y^{(j)}\perp{Z}$. Take the trial described in Figure 2 as an example, where $\pi_{j}({Z})$ varies with ${Z}$ (which does not occur in traditional clinical trials), the condition $Y^{(j)}\perp{Z}$ unrealistically rules out the influence of disease subtype on the outcome, as well as potential temporal effects, such as later recruited individuals being healthier than those recruited earlier. The problem with the naive estimator $\bar{Y}_{jk}$ arises from confounding by $\pi_{j}({Z})$, not from conditioning on $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$.
@@ 4.3 Inverse probability weighting
As noted in Section 4.2, the naive approach is biased because it does not account for confounding by $Z$. Under Assumption 1, since the true probability of receiving treatment, $\pi_{j}(Z)$, is known, this confounding can be addressed through inverse probability weighting (IPW). Notably, weighting by $\pi_{j}(Z)$ also helps correct for sample bias, as IPW uses only those assigned to treatment $j$, which might otherwise be a biased sample from the ECE population in platform trials (see the discussion at the end of Section 3.2).
The IPW identification formula is (the proof is in the Supplement):
$\theta_{jk}=E\bigg{\{}\frac{I(A=j)Y}{\pi_{j}({Z})}\ \Big{|}\ \pi_{j}({Z})>0,\pi_{k}({Z})>0\bigg{\}},$ (3)
@@
$\hat{\theta}_{jk}^{\rm\,ipw}=\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)Y_{i}}{\pi_{j}({Z}_{i})},$
@@
This IPW estimator may be improved by the following stabilized IPW (SIPW) (Hájek,, 1971) estimator, which normalizes the weights $I(A_{i}=j)/\pi_{j}({Z}_{i})$ for $i\in{{{\cal I}}}_{jk}$ to sum to 1:
@@
The IPW and SIPW methods utilize some covariate information through weighting with $\pi_{j}({Z})$. To improve efficiency, we can adjust for ${Z}$ and other baseline covariates using a model-assisted approach (Tsiatis et al.,, 2008; Ye et al.,, 2023; Bannick et al.,, 2025), which incorporates these covariates through a working model between the outcome and covariates and produces an asymptotically unbiased estimator even if the working model is incorrect. This robustness makes model-assisted approaches increasingly popular and recommended by the regulatory agencies (FDA, 2023a, ).
Specifically, let $\bm{X}\subset\bm{W}$ denote the observed vector of baseline covariates for adjustment, which may overlap with ${Z}$. We estimate $E(Y^{(j)}\mid\bm{X},\pi_{j}({Z})>0,\pi_{k}({Z})>0,A=j)$ by $\hat{\mu}_{jk}(\bm{X})$, using a working model that may be misspecified. For continuous outcomes, $\hat{\mu}_{jk}(\bm{X})$ is usually from least squares regression of $Y^{(j)}$ on $\bm{X}$ using data with $\pi_{j}({Z})>0,\pi_{k}({Z})>0$ and $A=j$. For binary outcomes, $\hat{\mu}_{jk}(\bm{X})$ is often obtained from a logistic working model. After $\hat{\mu}_{jk}(\cdot)$ is obtained, we can adjust for $\bm{X}$ using the augmented inverse probability weighting (AIPW) (Robins et al.,, 1994), resulting in the following AIPW estimator of $\theta_{jk}$,
@@
Interestingly, in its last term, the AIPW estimator leverages covariate information from all individuals within ${{{\cal I}}}_{jk}$, including those not receiving treatments $j$ or $k$. For example, in Figure 2(a), when estimating the effect of treatment 3 versus treatment 1 at the end of enrollment window 2, the AIPW also incorporates covariate data from subtype A individuals assigned to treatment 2 (in sub-study 1) and treatment 4 (in sub-study 3) during enrollment windows 1 and 2. These covariate data are not used in the IPW or SIPW estimators. Thus, AIPW offers a robust method to pool data from all individuals within ${{{\cal I}}}_{jk}$, thereby further extracting the potential of efficiency gain through platform trials.
@@
Both AIPW and SAIPW estimators are asymptotically normal (Theorem 1 in Section 5). Note that when $Z_{i}$ is discrete, $\pi_{j}(Z_{i})$ can be estimated using sample proportions and used in place of the true value $\pi_{j}(Z)$ in any of the above weighting estimators.